Imagine that the shot run price elasticity of supply for a framer's corn is 0.3, while the long run price elasticity is 2. If price for corn fall 30 percent, what are the short run and long run changes in quantity supplies? What are the short run and long run changes in quantity supplies if prices rise by 15 percent? what happens to the farmer's revenues in each of these situations?
1
Expert's answer
2013-03-13T09:28:19-0400
Imagine that the shot run price elasticity of supply for a framer's corn is
0.3, while the long run price elasticity is 2. If price for corn fall 30
percent, what are the short run and long run changes in quantity supplies?
What are the short run and long run changes in quantity supplies if prices
rise by 15 percent? what happens to the farmer's revenues in each of these
situations?
For 30% fall in price:
1) in the short run: change in quantity = 0.3*(-30%) = -9%, revenue fall, but
not sharply, as the supply is inelastic.
2) in the long run: change in quantity = 2*(-30%) = -60%, revenue fall sharply,
as the supply is elastic.
For 15% rise in price:
1) in the short run: change in quantity = 0.3*(+15%) = +4.5%, revenue rise, but
not sharply, as the supply is inelastic.
2) in the long run: change in quantity = 2*(+15%) = +30%, revenue rise sharply,
as the supply is elastic.